THE SUBMILLIMETER-WAVE SPECTRUM OF ALLYL ALCOHOL (CH_{2}=CHCH_{2}OH)

\underline{BRIAN J. DROUIN}, JOHN C. PEARSON, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099.

Two isomers of allyl alcohol have been further studied in selected regions of the submillimeter and microwave spectra. The skew-gauche and syn-gauche conformations have been previously identified at room temperature^{a}^{,}^{b} and in a free-jet^{c}, respectively. The present study positively identifies the syn-gauche conformer in the gas-phase at room temperature and confirms that tunneling, as opposed to internal rotation, of the hydroxyl proton is the predominant perturbation up to K_{a} = 30. Further measurements made in the X-band, at kHz precision, have increased the precision of the Coriolis coupling terms and energy level (tunneling) splitting. Extension of the spectral measurements has provided a set of spectral parameters that can accurately predict spectral line frequencies through much of the sub-millimeter.

aA. N. Murty and R. F. Curl, Jr., J. Chem. Phys. 46, 4176-4180 (1967)

bH. Badawi, P. Lorencak, K. W. Hillig II, M. Imachi and R. L. Kuczkowski, J. Mol. Struc. 162, 247-254 (1987).

cS. Melandri, P. G. Favero and W. Caminati, Chem. Phys. Let. 223, 541-545, (1994)